OpenStudy (anonymous):

Write an equation in slope-intercept form for the line passing through (5, 1) and (-4, 7)

OpenStudy (anonymous):

y-y1=m(x-x1)

Parth (parthkohli):

Standard form is y = mx + b First determine the slope. \(\Large \color{Black}{\Rightarrow {Slope = {y_2 - y_1 \over x_2 - x_1}} }\) Then substitute m with the slope and find b.

OpenStudy (anonymous):

where (x1,y1) are a point in the line. You are given two

OpenStudy (anonymous):

I agree with ParthKohli i am new so he typed it in faster XD

OpenStudy (anonymous):

choose 1 and then find the slope (m) with the given equation from Parthkohli

OpenStudy (anonymous):

But it's better to do it from point slope form and then isolate y. slope-intercept form is sometimes hard to see right away and you are not given the y-int as one of the points

OpenStudy (anonymous):

Option A: y=-2/3x+13/3 Option B: y=-2/3x-13/3 Option C: y=2/3x-13/3 Option D: y=2/3x+13/3

OpenStudy (anonymous):

Umm you might want give it a try first....we already told you how to do it. Don't be lazy

OpenStudy (anonymous):

I tried and got the first answer

OpenStudy (anonymous):

is the first answer correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

thank you!